### v1

@@ VI.1 Electromagnetic interactions in the non-relativistic limit
Thus, applying $H$ on $\psi\left(\mathbf{r},t\right)$, we obtain
@@
$\left(\boldsymbol{\Pi}\times\boldsymbol{\Pi}\right)\varphi={}$ $\left(\boldsymbol{p}-\frac{e}{c}\mathbf{A}\right)\times\left(\boldsymbol{p}-\frac{e}{c}\mathbf{A}\right)\varphi,$
[…]
@@ VI.2 Foldy-Wouthuysen transformation
$\beta\mathcal{O}={}$ $\mathcal{-O}\beta,$
@@
Previously, we noted that our goal is to eliminate the odd component from the Hamiltonian, as it links the upper and lower sections of the bispinor. To achieve this, we apply a transformation in the Hamiltonian, ensuring that in the resulting form the odd component is no longer present. To do this, we define the operator
@@
$H_{\Phi}={}$ $\beta mc^{2}+\beta\frac{1}{2m}\left(\boldsymbol{p}-\frac{e}{c}\mathbf{A}\right)^{2}-\frac{e\mathrm{\hbar}}{2mc}\beta\left(\boldsymbol{\Sigma}\cdot\mathbf{B}\right)-\frac{\beta\,\boldsymbol{p}^{4}}{8m^{3}c^{6}}+e\phi$
$-\frac{e\hbar^{2}}{8m^{2}c^{2}}\boldsymbol{\nabla}\cdot\mathbf{E}-\frac{ie\hbar^{2}}{8m^{2}c^{2}}\boldsymbol{\Sigma}\cdot(\boldsymbol{\nabla}\times\mathbf{E})-\frac{e\hbar}{4m^{2}c^{2}}\boldsymbol{\Sigma}\cdot(\mathbf{E}\times\boldsymbol{p}).$ (149)
@@
$H_{\Phi}={}$ $mc^{2}+\frac{1}{2m}\left(\boldsymbol{p}-\frac{e}{c}\mathbf{A}\right)^{2}-\frac{\boldsymbol{p}^{4}}{8m^{3}c^{6}}-\frac{e\mathrm{\hbar}}{2mc}\boldsymbol{\sigma}\cdot\mathbf{B}+e\phi$
$-\frac{ie\hbar^{2}}{8m^{2}c^{2}}\boldsymbol{\sigma}\cdot\left(\boldsymbol{\nabla}\times\mathbf{E}\right)-\frac{e\hbar}{4m^{2}c^{2}}\boldsymbol{\sigma}\cdot\left(\mathbf{E}\times\boldsymbol{p}\right)$
@@
The first three terms above represent the expansion of the kinetic energy $K=[\left(\boldsymbol{p}-e/c\mathbf{A}\right)^{2}+mc^{2}]^{1/2}-mc^{2}$, describing the increase in mass in the relativistic case. The next pair of terms is associated with magnetic dipole and electrostatic energy. Subsequently, the next two terms describe the spin-orbit (SO) interaction. For a spherically symmetric potential, we have ×$\nabla\times\mathbf{E}=0$. Additionally, we can write
$\boldsymbol{\sigma}\cdot(\boldsymbol{E}\times\boldsymbol{p})=-\frac{1}{r}\frac{\partial\phi}{\partial r}\boldsymbol{\sigma}\cdot(\mathbf{r}\times\boldsymbol{p})=-\frac{1}{r}\frac{\partial\phi}{\partial r}(\boldsymbol{\sigma}\cdot\mathbf{L}),$ (152)
@@ VII Relativistic covariant notation
$\left(-i\boldsymbol{\alpha}\cdot\boldsymbol{\nabla}+\beta m\right)\psi\left(\mathbf{r},t\right)=i\frac{\partial\psi\left(\mathbf{r},t\right)}{\partial t}.$ (155)
@@
$\frac{\partial}{\partial x^{\mu}}={}$ $=\left(\frac{\partial}{\partial x^{0}},\frac{\partial}{\partial x^{1}},\frac{\partial}{\partial x^{2}},\frac{\partial}{\partial x^{3}}\right),$
[…]
@@
$p_{\mu}={}$ $\left(p^{0},-\boldsymbol{p}\right),$ (169)
$p_{\mu}={}$ $\left(E,-\boldsymbol{p}\right),$
@@
$p_{\mu}p^{\mu}={}$ $\left(p^{0}\right)^{2}-\boldsymbol{p}^{2},$
[…]
@@
must be invariant [89, 5]. Consequently, it can be inferred that the combination of derivatives, expressed in covariant notation as $\partial_{\mu}\partial^{\mu}$, remains constant when transitioning between inertial frames of reference.
@@
[…]
@@
[…]
@@
Furthermore, if one wishes to write the Dirac equation even more bluntly, Feynman notation, where $\gamma^{\mu}\partial_{\mu}=\not{\partial}$̸, is employed, so that
@@
Feature Schrödinger Klein-Gordon Dirac Equation $i\hbar\frac{\partial\psi}{\partial t}=\left(-\frac{\hbar^{2}}{2m}\boldsymbol{\nabla}^{2}+V\right)\psi$ $\frac{1}{c^{2}}\frac{\partial^{2}\psi}{\partial t^{2}}=\left(\boldsymbol{\nabla}^{2}-\left(\frac{mc}{\hbar}\right)^{2}\right)\psi$ $i\hbar\frac{\partial\psi}{\partial t}=\left(-ic\hbar\boldsymbol{\alpha}\cdot\boldsymbol{\nabla}+\beta mc^{2}\right)\psi$ Energy Only positive eigenvalues. Both positive and negative energies are allowed. Both positive and negative energies are allowed. Probabilistic interpretation Viable, since $\rho>0$. Not viable since $\rho$ is not positive definite. Viable, since $\rho>0$. Relativity Not-invariant under Lorentz transformation. Invariant under Lorentz transformation. Invariant under Lorentz transformation. Spin Does not contain spin. Does not contain spin. Contains the spin.
@@ VIII Conclusions
We started with a historical preamble, illustrating that, in the 1920s: i) Physics was effervescent, marked by new discoveries in quantum mechanics, ii) a consensus on a consistent methodology had not yet been reached, and iii) Dirac was deeply involved in both previous aspects, actively contributing to the construction of the new physics. Moreover, we outlined features of matrix mechanics and underscored the significance of Dirac’s canonical quantization, which was found to be applicable in other instances in this study.
The natural course was to present wave mechanics as an alternative to matrix mechanics. Schrödinger’s equation enabled many physicists [32] to have a more practical view of quantum mechanics and, from it, Born was able to extract a probabilistic interpretation that originated from the density $\rho(\mathbf{r},t)$ being positive definite. Furthermore, Schrödinger managed to solve the problem of the eigenvalues for the hydrogen atom, deriving the non-relativistic hydrogen spectra – a feat that matrix mechanics had not been able to accomplish without ad hoc hypotheses. However, Schrödinger equation is not invariant under the Lorentz transformation; it does not fit into relativity. We show its non-invariance explicitly, with a one-dimensional Lorentz transformation and without relativistic notation.
@@
We encountered negative energy solutions once again after solving the dynamics for a free particle at rest. The deeper investigation into the physical significance of these energies unfolds when considering the case of a moving particle, where we have observed that the two upper components of the bispinor correspond to the particle, associated with positive energy, while the lower components correspond to an entity linked with negative energy. We commented on how Dirac’s conjecture, his sea of negative energy, explains this result and postulates, through a “hole” in the sea, the existence of an antiparticle. Helicity serves as an auxiliary means of interpreting these outcomes: Positive (right-handed) and negative (left-handed) helicity are imposed, respectively, on the first and second components of the upper part of the bispinor in the case of positive energy. For negative energy, the third component of the bispinor exhibits right-hand helicity, while the fourth component displays left-hand helicity. We have used these types of helicity to discern the spin of the solutions that emerge from the free-particle problem, just as we employed the sign of energy to differentiate between the particle and the antiparticle.

### v2

@@ VI.1 Electromagnetic interactions in the non-relativistic limit
Thus, applying $H$ on $\Psi\left(\mathbf{r},t\right)$, Eq. (91), we obtain
@@
$\left(\boldsymbol{\Pi}\times\boldsymbol{\Pi}\right)\varphi={}$ $\left(\mathbf{p}-\frac{e}{c}\mathbf{A}\right)\times\left(\mathbf{p}-\frac{e}{c}\mathbf{A}\right)\varphi,$
$={}$ $\left(-i\hbar\boldsymbol{\nabla}-\frac{e}{c}\mathbf{A}\right)\times\left(-i\hbar\boldsymbol{\nabla}\varphi-\frac{e}{c}\mathbf{A}\varphi\right),$
$={}$ $i\hbar\frac{e}{c}\left[\boldsymbol{\nabla}\times\left(\mathbf{A}\varphi\right)+\mathbf{A}\times\left(\boldsymbol{\nabla}\varphi\right)\right],$
$={}$ $i\hbar\frac{e}{c}\left(\boldsymbol{\nabla}\times\mathbf{A}\right)\varphi,$
$={}$ $i\hbar\frac{e}{c}\mathbf{B}\,\varphi,$ (129)
@@ VI.2 Foldy-Wouthuysen transformation
$\beta\,\mathcal{O}={}$ $-\mathcal{O}\,\beta,$
@@
Previously, we noted that our goal is to eliminate the odd component from the Hamiltonian, since it links the upper and lower sections of the bispinor. To achieve this, we apply a transformation in the Hamiltonian, ensuring that in the resulting form the odd component is no longer present. To do this, we define the operator
@@
$H_{\Phi}={}$ $\beta mc^{2}+\beta\frac{1}{2m}\left(\mathbf{p}-\frac{e}{c}\mathbf{A}\right)^{2}-\frac{e\mathrm{\hbar}}{2mc}\beta\left(\boldsymbol{\Sigma}\cdot\mathbf{B}\right)-\frac{\beta\,\mathbf{p}^{4}}{8m^{3}c^{6}}+e\phi$
$-\frac{e\hbar^{2}}{8m^{2}c^{2}}\boldsymbol{\nabla}\cdot\mathbf{E}-\frac{ie\hbar^{2}}{8m^{2}c^{2}}\boldsymbol{\Sigma}\cdot(\boldsymbol{\nabla}\times\mathbf{E})-\frac{e\hbar}{4m^{2}c^{2}}\boldsymbol{\Sigma}\cdot(\mathbf{E}\times\mathbf{p}).$ (150)
@@
$H_{\Phi}={}$ $mc^{2}+\frac{1}{2m}\left(\mathbf{p}-\frac{e}{c}\mathbf{A}\right)^{2}-\frac{\mathbf{p}^{4}}{8m^{3}c^{6}}-\frac{e\mathrm{\hbar}}{2mc}\boldsymbol{\sigma}\cdot\mathbf{B}+e\phi$
$-\frac{ie\hbar^{2}}{8m^{2}c^{2}}\boldsymbol{\sigma}\cdot\left(\boldsymbol{\nabla}\times\mathbf{E}\right)-\frac{e\hbar}{4m^{2}c^{2}}\boldsymbol{\sigma}\cdot\left(\mathbf{E}\times\mathbf{p}\right)$
@@
The first three terms above represent the expansion of the kinetic energy $K=[\left(\mathbf{p}-e/c\mathbf{A}\right)^{2}+mc^{2}]^{1/2}-mc^{2}$, describing the increase in mass in the relativistic case. The next pair of terms is associated with magnetic dipole and electrostatic energy. Subsequently, the next two terms describe the spin-orbit (SO) interaction. For a spherically symmetric potential, we have $\nabla\times\mathbf{E}=0$. Additionally, we can write
$\boldsymbol{\sigma}\cdot(\boldsymbol{E}\times\mathbf{p})=-\frac{1}{r}\frac{\partial\phi}{\partial r}\boldsymbol{\sigma}\cdot(\mathbf{r}\times\mathbf{p})=-\frac{1}{r}\frac{\partial\phi}{\partial r}(\boldsymbol{\sigma}\cdot\mathbf{L}),$ (153)
@@
There are many other interesting phenomena related to the Dirac equation. In this context, we cite two topics that highlight novel aspects of the equation and serve as extensions of the themes discussed here. Interested readers may further explore them by treating these topics as exercises, with guidance from appropriate textbooks [16, 62]. These include:
  1. 1.
    Study the Zitterbewegung effect by computing the Heisenberg equation of motion for the position operator.
  2. 2.
    Study the Klein paradox, by deriving the transmission and reflection coefficients in the scattering of an electron hitting a potential step [70].
@@ VII Relativistic covariant notation
$\left(-i\boldsymbol{\alpha}\cdot\boldsymbol{\nabla}+\beta m\right)\psi\left(\mathbf{r},t\right)={}$ $i\frac{\partial}{\partial t}\psi\left(\mathbf{r},t\right).$ (156)
@@
$\frac{\partial}{\partial x^{\mu}}={}$ $\left(\frac{\partial}{\partial x^{0}},\frac{\partial}{\partial x^{1}},\frac{\partial}{\partial x^{2}},\frac{\partial}{\partial x^{3}}\right),$
$={}$ $\left(\partial_{0},\partial_{1},\partial_{2},\partial_{3}\right),$
$={}$ $(\partial_{t},\boldsymbol{\nabla}),$
$={}$ $\partial_{\mu},$ (167)
@@
$p_{\mu}={}$ $\left(p^{0},-\mathbf{p}\right),$ (170)
$p_{\mu}={}$ $\left(E,-\mathbf{p}\right),$
@@
$p_{\mu}p^{\mu}={}$ $\left(p^{0}\right)^{2}-\mathbf{p}^{2},$
$={}$ $E^{2}-\mathbf{p}^{2},$
$={}$ $m^{2},$ (171)
@@
must be invariant [87]. Consequently, it can be inferred that the combination of derivatives, expressed in covariant notation as $\partial_{\mu}\partial^{\mu}$, remains constant when transitioning between inertial frames of reference.
@@
$={}$ $\gamma\left(a^{0}-\beta a^{1}\right)\gamma\left(b_{0}-\beta b_{1}\right)$
@@
$={}$ $\gamma^{2}\left(1-\beta^{2}\right)a^{0}b_{0}-\gamma^{2}\left(1-\beta^{2}\right)a^{1}b_{1}-a^{2}b^{2}-a^{3}b^{3},$ (176)
@@
Furthermore, if one wishes to write the Dirac equation even more bluntly, Feynman notation, where $\gamma^{\mu}\partial_{\mu}=\not{\partial}$, is employed, so that
@@
Feature Schrödinger Klein-Gordon Dirac Equation $i\hbar\frac{\partial}{\partial t}\psi=\left(-\frac{\hbar^{2}}{2m}\boldsymbol{\nabla}^{2}+V\right)\psi$ $\frac{1}{c^{2}}\frac{\partial^{2}}{\partial t^{2}}\psi=\left(\boldsymbol{\nabla}^{2}-\left(\frac{mc}{\hbar}\right)^{2}\right)\psi$ $i\hbar\frac{\partial}{\partial t}\psi=\left(-ic\hbar\boldsymbol{\alpha}\cdot\boldsymbol{\nabla}+\beta mc^{2}\right)\psi$ Energy Only positive eigenvalues. Both positive and negative energies are allowed. Both positive and negative energies are allowed. Probabilistic interpretation Viable, since $\rho>0$. Not viable since $\rho$ is not positive definite. Viable, since $\rho>0$. Relativity Not-invariant under Lorentz transformation. Invariant under Lorentz transformation. Invariant under Lorentz transformation. Spin Does not contain spin. Does not contain spin. Contains the spin.
@@ VIII Summary
We started with a historical preamble, illustrating that, in the 1920s: i) Physics was in a period of rapid development, marked by new discoveries in Quantum Mechanics, ii) a consensus on a consistent methodology had not yet been reached, and iii) Dirac was deeply involved in both previous aspects, actively contributing to the construction of the new physics. Moreover, we outlined features of matrix mechanics and underscored the significance of Dirac’s canonical quantization, which was found to be applicable in other instances in this study.
The natural course was to present wave mechanics as an alternative to matrix mechanics. Schrödinger’s equation enabled many physicists [31] to have a more practical view of Quantum Mechanics and, from it, Born was able to extract a probabilistic interpretation that originated from the density $\rho(\mathbf{r},t)$ being positive definite. Furthermore, Schrödinger managed to solve the problem of the eigenvalues for the hydrogen atom, deriving the non-relativistic hydrogen spectra – a feat that matrix mechanics had not been able to accomplish without ad hoc hypotheses. However, the Schrödinger equation is not invariant under the Lorentz transformation; it does not fit into Relativity. We show its non-invariance explicitly, with a one-dimensional Lorentz transformation and without relativistic notation.
@@
We encountered negative energy solutions once again after solving the dynamics for a free particle at rest. The deeper investigation into the physical significance of these energies unfolds when considering the case of a moving particle, where we have observed that the two upper components of the bispinor correspond to the particle, associated with positive energy, while the lower components correspond to an entity linked with negative energy. We commented on how Dirac’s conjecture, his sea of negative energy, explains this result and postulates, through a “hole” in the sea, the existence of an antiparticle. Helicity serves as an auxiliary means of interpreting these outcomes: Positive (right-handed) and negative (left-handed) helicity are imposed, respectively, on the first and second components of the upper part of the bispinor in the case of positive energy. For negative energy, the third component of the bispinor exhibits right-hand helicity, while the fourth component displays left-hand helicity. We have used these types of helicity to discern the spin of the solutions that emerge from the free particle problem, just as we employed the sign of energy to differentiate between the particle and the antiparticle.